a father in now 38 years older than his son ten years ago he was twice as older as his son how old are the father and son now?
Answer by Cromlix(4381)

Hi there,
a father in now 38 years older than his son
F = 38 + S..........(1)
ten years ago he was twice as older as his son
F - 10 = 2(S - 10)........(2)
Replace F in Eq(2) with 38 + S from Eq (1)
38 + S - 10 = 2(S - 10)
38 + S - 10 = 2S - 20
Collect like terms
38 - 10 + 20 = 2S - S
48 = S
Son is 48 years old
Father is 86 years old.
Hope this helps :-)